# Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series

### v1

@@ 4.2 General missing pattern: A two-state Markov chain example
In this subsection, the identification issue of the temporal Wasserstein imputation method is studied via a two-state Markov chain example. Particularly, we connect the non-identification of the underlying marginal distributions to the lack of uniqueness of the solutions to a system of linear equations. We also show that, under two favorable conditions, the correct marginal distributions can still be pinned down by TWI. To free our discussion from finite-sample complications, we will focus on the distribution level in this subsection, or intuitively speaking, assume “infinite sample.”
@@
$\mathbb{P}(\tilde{x}_{t}=1|\tilde{x}_{t-1}=0)=0.2,$
$\mathbb{P}(\tilde{x}_{t}=0|\tilde{x}_{t-1}=1)=0.6.$
The transition probabilities are specified to facilitate discussions; the main results below are not specific to these choices. It is easy to check that $\{\tilde{x}_{t}\}$ has the following stationary distribution.
$\mathbb{P}(\tilde{x}_{t}=1)=0.75,\quad\mathbb{P}(\tilde{x}_{t}=0)=0.25.$
Consequently, the 2-dimensional marginal distribution of $\{\tilde{x}_{t}\}$ is specified by
@@
Without loss of generality, we set $n_{1}=0$. Suppose that before time $n_{1}$, there is a missing value every $q_{1}$ observations. Likewise, there is a missing value every $q_{2}$ observations after time $n_{1}$. Specifically, $x_{t}=\bigstar$ if $-t$ is a positive integer multiple of $q_{1}$, and similarly, $x_{t}=\bigstar$ if $t$ is a positive integer multiple of $q_{2}$, where $q_{1}$, $q_{2}$ are some integers larger than 2. Next, suppose an imputation $\{\hat{w}_{t}\}_{-\infty<t<\infty}$ has been made such that if $t<n_{1}$,
@@
for some $a_{1},a_{2},b_{1},b_{2}\in[0,1]$. Note that $\mathbb{P}$ in (14) and (15) should be interpreted as proportions since the imputation may still depend on the whole sequence of $\{x_{t}\}$, and it is assumed that they are well-defined. Imputation methods corresponding to varying $a_{1},a_{2},b_{1},b_{2}$ will lead to different 2-dimensional marginal distributions. Indeed, from (13)–(15), we can compute the 2-dimensional marginal distributions of the imputed series. For $\{\hat{w}_{t}\}_{-\infty<t<0}$,
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1))$ $=0.6(1-q_{1}^{-1})+\frac{3}{4q_{1}}a_{1},$
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0))$ $=0.15(1-q_{1}^{-1})+\frac{1}{4q_{1}}b_{1},$
@@
Clearly, if $q_{1}=q_{2}$, then any imputation with $a_{1}=a_{2}$ and $b_{1}=b_{2}$ satisfies the above system of equations. Therefore, identification is not possible in this case, which leads to our first condition.
Asymmetric missing pattern: $q_{1}\neq q_{2}$.
However, condition (C1) is still inadequate for identification because the equations in (17) are not linearly independent. The lack of uniqueness of the solutions to (17) is at the crux of the identification issue of the imputation problem. The blue lines in Figure 1 plots the set of solutions to (17). Each point on the blue line represents a legitimate TWI imputation whose marginal distributions before and after time $n_{1}$ are the same.
Nonetheless, (16) hints an additional requirement for the correct imputation, giving rise to the following condition.
@@
That is, the only marginal distribution that achieves zero Wasserstein loss and satisfies (C2) is that of the underlying distribution.
As shown in Figure 1, the 45 degree line (dashed orange), which imposes (C2), intersects with the set of temporal Wasserstein imputations at the correct calibrations.
The above example illustrates when TWI can potentially identify the underlying marginal distribution. The crucial interpolator stability condition (C2) thus provides some guidance on the selection of TWI solutions. Indeed, Algorithm 1 may converge to varying solutions depending on the initialization. (C2) implies that we shall choose the solution so that the relationship between the imputed values and the neighboring observations stays stable before and after $n_{1}$. In higher dimensions, summary statistics can be used. For example, fit a linear regression of imputed values on neighboring values using data before and after $n_{1}$ and compare the similarity of the coefficients. To promote interpolator stability, we suggest running TWI with different cut-off $n_{1}$, using the previously obtained imputations as initializations. This leads to a simple procedure termed $k$-TWI henceforth, which is summarized in Algorithm 2 below.
@@ 5 Simulation Studies
In this section, we apply the proposed temporal Wasserstein imputation to data generated from various time series models, including multivariate, nonlinear, and nonstationary ones, to examine its performance. For comparison, we employ alternative imputation methods such as Kalman smoothing, the scalar filtering algorithm of Peña and Tsay, (2021), and deterministic methods such as linear and spline interpolations. In addition, the following two missing patterns are considered.
  • Pattern I: 300 observations are omitted at random, which corresponds to sporadic missing data.
  • Pattern II: For every 20 observations, 6 consecutive observations are omitted, which mimics patches of missing data.
The total length of the time series is set to $n=1000$. Thus, for both missing patterns, 30% of the data are missing. For each model and missing pattern considered in this section, 1000 Monte Carlo simulations are carried out to compute the measure of performance.
We employ the proximal gradient descent (Parikh et al.,, 2014) to efficiently solve step (b) in Algorithm 1 when no side information is available since the projection step is straightforward. Conversely, when $\mathcal{C}$ is defined by a system of linear equations, we solve step (b) using the closed form (7). In addition, we apply the $k$-TWI (Algorithm 2) with three cut-off points ($0.25n$, $0.5n$, and $0.75n$) to the simulated data. Both linear interpolation and Kalman smoothing are used as initialization of TWI in our experiments. Kalman smoothing and the deterministic smoothing techniques are implemented via the imputeTS package (Moritz and Bartz-Beielstein,, 2017) in R. The scalar filter algorithm (denoted as ScalarF henceforth) utilizes intervention analysis, and can be applied to multivariate and nonstationary series. The details are documented in Appendix B.
@@ 5.1 Linear and nonlinear univariate time series
In this subsection, we focus on univariate time series exhibiting linear or nonlinear dynamic patterns. Specifically, we consider the following time series models. In the sequel, $\{\epsilon_{t}\}$ and $\{\epsilon_{t,j}\}$ for $j=1,2,\ldots$ denote i.i.d. sequences of standard Gaussian variables.
Model 1 (AR).
@@
Model 4 (I(1) process).
Data $\{x_{t}\}$ are generated according to
$z_{t}=$ $\phi z_{t-1}+0.5\epsilon_{t,1},$
$x_{t}=$ $x_{t-1}+z_{t}+\epsilon_{t,2},$
with $\phi=-0.7$.
Model 5 (Deterministic cyclic series with noise (CYC)).
@@
Models 1 and 2 are linear time series models commonly seen in practice. Kalman smoothing is particularly adept at imputing such series. Note that Model 2 admits an AR($\infty$) representation, so any finite-dimensional marginal distribution, the key objects TWI and $k$-TWI work with, is inadequate to fully characterize the distribution of the underlying process. Model 3 is a celebrated threshold AR model which has found many applications in modeling nonlinear dynamics in the econometrics and statistics literature (Montgomery et al.,, 1998; Li and Ling,, 2012; Tsay and Chen,, 2018). Model 4 generates an I(1) series which typically produces a highly persistent stochastic trend. To remove the unit root, difference transform is usually applied to the data $\{x_{t}\}$; that is, one often deals with $\nabla x_{t}=x_{t}-x_{t-1}$ instead of the original series. However, this may result in excessive missing values in $\{\nabla x_{t}\}$, which can cause difficulty in imputation. Finally, Model 5 features a deterministic cyclic trend with non-standard periods. Specifically, it includes two sinusoidal components with periods of 8.70 and 11.77, respectively.
@@
as a measure of how well the imputed series approximates the original series. Second, we compare the root mean square estimation errors (RMSE) of the estimated autocovariance functions. Define $\gamma(\mathbf{z},h)=(n-h)^{-1}\sum_{t=h}^{n-1}(z_{t}-\bar{z})(z_{t-h}-\bar{z})$, where $\mathbf{z}=(z_{0},z_{1},\ldots,z_{n-1})$ and $\bar{z}=n^{-1}\sum_{t=0}^{n-1}z_{t}$. Then the RMSE is defined as
$\mathrm{RMSE}_{\gamma}(h)=\sqrt{\frac{1}{1000}\sum_{i=1}^{1000}(\gamma(\hat{\mathbf{w}}^{(i)},h)-\gamma_{\mathbf{x}}(h))^{2}},$ (19)
where $\hat{\mathbf{w}}^{(i)}=(\hat{w}_{0}^{(i)},\ldots,\hat{w}_{n-1}^{(i)})$ and $\gamma_{\mathbf{x}}(h)=\frac{1}{1000}\sum_{i=1}^{1000}\gamma(\mathbf{x}^{(i)},h)$ with $\mathbf{x}^{(i)}=(x_{0}^{(i)},\ldots,x_{n-1}^{(i)})$. Third, for Models 14, the model parameters are estimated using the imputed series, and their root mean square errors are reported.
@@
The Wasserstein distance between the imputed and the original marginal distributions is reported in Table 1. Across all five models employed, the smallest loss is consistently achieved by TWI-type methods, indicating that TWI effectively approximates the distributional properties of the underlying series. While other methods may yield good imputations for some specific models, the result is not consistent across all models. For instance, linear and spline interpolations show low Wasserstein loss for Model 1, but their performance is less competitive when applied to other models. Similarly, the Kalman smoothing performs well under Model 1 with missing pattern I, but its advantage diminishes under missing pattern II. In contrast, TWI methods consistently deliver solid performance across models and missing patterns, provided they are carefully initialized.
@@
Additionally, Table 1 also shows TWI typically improves the Wasserstein loss from its initialization, often siginificantly. For example, for Model 3 (TAR), TWI reduces the Wasserstein loss from Kalman smoothing initialization by approximately 45% under missing pattern I (from 1.38 to 0.74). Similarly, for Model 5 (CYC), TWI reduces the Wasserstein loss by more than 50% from linear initialization (from 1.96 to 0.79). Figure 2 plots $\hat{w}_{t}$ against $\hat{w}_{t-1}$ when either or both are imputed (${x}_{t}=\bigstar$ or ${x}_{t-1}=\bigstar$) under Model 3, alongside $(x_{t},x_{t-1})$ of the original data. While Kalman smoothing introduces substantial distortion in the marginal distribution, TWI, initialized by Kalman smoothing, significantly corrects this bias, producing imputations that capture the nonlinearity in the ground truth. Figure 3 similarly plots $\hat{w}_{t}$ against $\hat{w}_{t-h}$ (and $x_{t}$ against $x_{t-h}$) for $h=1,2,6,10$ under Model 5. As shown in the figure, TWI markedly improves from its linear initialization by “denoising” the marginal distributions, leading to a distribution that closely match the ground truth.
For the nonstationary Model 4 (I(1)), we apply the methods to obtain an imputation $\{\hat{w}_{t}\}$ and focus on the differenced series $\nabla\hat{w}_{t}=\hat{w}_{t}-\hat{w}_{t-1}$. Table 1 shows the Wasserstein distance between the marginal distributions implied by $\{\nabla\hat{w}_{t}\}$ and $\{\nabla x_{t}\}$. As noted before, there are many more missing values in $\{\nabla x_{t}\}$ than originally in $\{x_{t}\}$. In this case, TWI still consistently yields the lowest Wasserstein loss and greatly improves from its initialization.
@@
Table 1: Wasserstein distance between the empirical marginal distributions of the imputed series and those of the full data, averaged over 1000 simulations. TWIlin denotes temporal Wasserstein imputation using linear interpolation as initialization and TWIKal denotes temporal Wasserstein imputation with Kalman smoothing as initialization. Similar notations are used for $k$-TWI. For Model 4 (I(1)), the loss is computed using the first-differenced series. The smallest value for each model and missing pattern is marked in boldface.
@@
ALR (×$\times 10$) 0.56 1.83 3.81 0.46 0.36 0.33 0.40 0.34
(a) Marginal distribution under missing pattern I
(b) Marginal distribution under missing pattern II
@@
(a) $h=1$
(b) $h=2$
(c) $h=6$
(d) $h=10$

### v2

@@ 4.2 General missing pattern: A two-state Markov chain example
In this subsection, the identification issue of the temporal Wasserstein imputation method is studied via a two-state Markov chain example. Particularly, we connect the non-identification of the underlying marginal distributions to the lack of uniqueness of the solutions to a system of linear equations. We also show that, under two natural conditions, the correct marginal distributions can still be pinned down by TWI. To free our discussion from finite-sample complications, we will focus on the distribution level in this subsection, or intuitively speaking, assume “infinite sample.”
@@
$\mathbb{P}(\tilde{x}_{t}=1|\tilde{x}_{t-1}=0)=0.6,\quad$ $\mathbb{P}(\tilde{x}_{t}=0|\tilde{x}_{t-1}=1)=0.2.$
The transition probabilities are specified to facilitate discussions, but the main results below are not specific to these choices. It is easy to check that $\{\tilde{x}_{t}\}$ has the stationary distribution given by $\mathbb{P}(\tilde{x}_{t}=1)=0.75$ and $\mathbb{P}(\tilde{x}_{t}=0)=0.25$. Consequently, the 2-dimensional marginal distribution of $\{\tilde{x}_{t}\}$ is specified by
@@
Without loss of generality, we set $n_{1}=0$. Suppose that before time $n_{1}$, there is a missing value every $q_{1}$ observations. Likewise, there is a missing value every $q_{2}$ observations after time $n_{1}$. Specifically, $x_{t}=\bigstar$ if $-t$ is a positive integer multiple of $q_{1}$, and similarly, $x_{t}=\bigstar$ if $t$ is a positive integer multiple of $q_{2}$, where $q_{1}$ and $q_{2}$ are some integers larger than 2 and $q_{1}\neq q_{2}$ (see condition (C1) below). Next, suppose an imputation $\{\hat{w}_{t}\}_{-\infty<t<\infty}$ has been made such that if $t<n_{1}$,
@@
for some $a_{1},a_{2},b_{1},b_{2}\in[0,1]$. Note that $\mathbb{P}$ in (10) and (11) should be interpreted as proportions since the imputation may depend on the whole sequence of $\{x_{t}\}$, and it is assumed that they are well-defined. Essentially, here we parametrize all imputation methods by $a_{1},a_{2},b_{1},b_{2}$, which correspond to different 2-dimensional marginal distributions. Indeed, from (9)–(11), we can compute the 2-dimensional marginal distributions of the imputed series. For $\{\hat{w}_{t}\}_{-\infty<t<0}$,
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1))$ $=0.6(1-\frac{1}{q_{1}})+\frac{3a_{1}}{4q_{1}},\quad\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0))=0.15(1-\frac{1}{q_{1}})+\frac{b_{1}}{4q_{1}},$
@@
If $q_{1}=q_{2}$, then any imputation with $a_{1}=a_{2}$ and $b_{1}=b_{2}$ would satisfy (13). Naturally, identification is not possible in this case, which leads to our first condition.
Non-systematic missing pattern: $q_{1}\neq q_{2}$.
Condition (C1) precludes the special case of down-sampling, which is notoriously hard to be treated as a missing data problem.
However, (C1) is inadequate for identification because the equations in (13) are not linearly independent. The lack of uniqueness of the solutions to (13) is at the crux of the identification problem in imputation. The blue lines in Figure 1 plots the set of solutions to (13). Each point on the blue line represents a legitimate TWI imputation whose marginal distributions before and after time $n_{1}$ are the same. Nonetheless, (12) hints an additional requirement for the correct imputation, giving rise to the following condition.
@@
That is, the only 2-dimensional marginal distribution that achieves zero Wasserstein loss and satisfies (C2) is that of the underlying distribution.
As shown in Figure 1, the 45 degree line (dashed orange), which imposes (C2), intersects with the solution set of temporal Wasserstein imputations at the correct calibrations. This example shows, under two natural requirements, TWI can achieve distributional identification through minimizing the Wasserstein distance between the marginal distributions.
In practice, appropriate tuning of $n_{1}$ and $p$ is needed for TWI to find an imputation with marginal distributions approximating the underlying ones. As shown in Proposition 5, it is not necessary to use an $n_{1}$ such that all missing data lie on the same side. Instead, we can select $n_{1}$ from a pre-specified interval $[t_{1},t_{2}]$ which yields the smallest Wasserstein loss, where $t_{2}-t_{1}$, $t_{2}$, and $n-t_{1}$ are sufficiently large. Alternatively, a simple procedure to promote interpolator stability, as suggested by (C2), is to run TWI with different cut-off $n_{1}$, using previously obtained imputations as initializations. This procedure, which we call $k$-TWI, is summarized in Algorithm 2. Finally, as a nonparametric method, $p$ can be chosen to diverge with the sample size with $p/n\rightarrow 0$.
@@ 5 Simulation Studies
In this section, we apply TWI to data generated from various time series models, including multivariate, nonlinear, and nonstationary ones, to examine its performance. For comparison, we employ imputation methods such as Kalman smoothing, the scalar filtering algorithm of Peña and Tsay, (2021), and deterministic methods such as linear and spline interpolations. In addition, the following two missing patterns are considered.
  • Pattern I: 300 observations are omitted at random;
  • Pattern II: For every 20 observations, 6 consecutive observations are omitted.
Pattern I simulates sporadic missing data while Pattern II generates patches of missing data. The total length of the time series is set to $n=1000$. Thus, for both missing patterns, 30% of the data are missing. For each data generating process and missing pattern considered, 1000 Monte Carlo simulations are carried out.
We employ the proximal gradient descent (Parikh et al.,, 2014) to efficiently solve step (b) in Algorithm 1 when no side information is available since the projection step is straightforward. Conversely, when $\mathcal{C}$ is defined by a system of linear equations, we solve step (b) using the closed form (6). Throughout we fix $n_{1}=0.4n$ and $p=6$ for TWI. In addition, we apply the $k$-TWI (Algorithm 2) with three cut-off points ($n_{1}\in\{0.25n,0.5n,0.75n\}$) to the simulated data. For the initializations of TWI, both linear interpolation and Kalman smoothing are used in our experiments. Kalman smoothing and the deterministic smoothing techniques are implemented via the imputeTS package (Moritz and Bartz-Beielstein,, 2017) in R. The scalar filter algorithm (denoted as ScalarF henceforth) utilizes intervention analysis, and can be applied to multivariate and nonstationary series. Its implementation details are documented in Section S2 of the supplementary material.
@@ 5.1 Linear and nonlinear univariate time series
In this subsection, we consider the following data-generating processes (DGPs) for univariate time series, which can exhibit linear or nonlinear dynamic patterns. In the sequel, $\{\epsilon_{t}\}$ and $\{\epsilon_{t,j}\}$ for $j=1,2,\ldots$ denote i.i.d. sequences of standard Gaussian variables.
DGP 1 (AR).
@@
DGP 4 (I(1) process).
Data $\{x_{t}\}$ are generated according to $x_{t}=x_{t-1}+z_{t}+\epsilon_{t,2}$, where $z_{t}=\phi z_{t-1}+0.5\epsilon_{t,1}$ with $\phi=-0.7$.
DGP 5 (Cyclic series with noise (CYC)).
@@
DGPs 1 and 2 are well-known linear time series models. Note that DGP 2 admits an AR($\infty$) representation, so any finite-dimensional marginal distribution is inadequate to fully characterize the distribution of the underlying process. DGP 3 is a celebrated threshold AR model which has found many applications in modeling nonlinear dynamics in the econometrics and statistics literature (Montgomery et al.,, 1998; Li and Ling,, 2012; Tsay and Chen,, 2018). DGP 4 generates an I(1) series which typically produces a highly persistent stochastic trend. Finally, DGP 5 features a deterministic cyclic trend with non-standard periods. Specifically, it includes two sinusoidal components with periods of 8.70 and 11.77, respectively.
@@
as a measure of how well the imputed series approximates the original series. Second, for DGP 14, the model parameters are estimated using the imputed series, and their root mean square errors (RMSEs) are reported. Third, we compare the RMSEs of the estimated autocovariance functions (ACFs). However, due to space constraints, we relegate the comparison of ACFs to Section S3.1 in the supplementary material.
@@
The Wasserstein distance between the imputed and the original marginal distributions is reported in Table 1. Across all five DGPs employed, the smallest loss is consistently achieved by TWI-type methods, indicating that TWI effectively approximates the distributional properties of the underlying series. While other methods may yield good imputations for some specific DGPs, the result is not consistent across all processes. For instance, linear and spline interpolations show low Wasserstein loss for DGP 1, but their performance is less competitive when applied to other models. In contrast, TWI methods, carefully initialized, consistently deliver solid performance across models and missing patterns. For the nonstationary Model 4 (I(1)), we first obtain an imputation $\{\hat{w}_{t}\}$ and consider the differenced series $\nabla\hat{w}_{t}=\hat{w}_{t}-\hat{w}_{t-1}$. Table 1 shows the Wasserstein distance between the marginal distributions implied by $\{\nabla\hat{w}_{t}\}$ and $\{\nabla x_{t}\}$. In this case, TWI consistently yields the lowest Wasserstein loss and greatly improves from its initialization.
@@
Table 1 also shows TWI typically improves the Wasserstein loss from its initialization, often significantly. For example, for DGP 3 (TAR), TWI reduces the Wasserstein loss from Kalman smoothing initialization by approximately 45% under missing pattern I (from 1.38 to 0.74). Similarly, for DGP 5 (CYC), TWI reduces the Wasserstein loss by more than 50% from linear initialization (from 1.96 to 0.79). Figure 2 plots $\hat{w}_{t}$ against $\hat{w}_{t-1}$ when either or both are imputed under DGP 3, alongside $(x_{t},x_{t-1})$ of the original data. While Kalman smoothing introduces substantial distortion in the marginal distribution, TWI, initialized by Kalman smoothing, significantly corrects this bias, producing imputations that capture the nonlinearity in the ground truth. Note that Kalman smoothing is done via an ARIMA state-space representation; hence the model is misspecified if the underlying process is nonlinear. Nevertheless, it already serves as a good initialization of TWI.
@@
Table 1: Wasserstein distance between the empirical marginal distributions of the imputed series and those of the full data, averaged over 1000 simulations. TWIlin denotes temporal Wasserstein imputation using linear interpolation as initialization and TWIKal denotes temporal Wasserstein imputation with Kalman smoothing as initialization. Similar notations are used for $k$-TWI. For DGP 4 (I(1)), the loss is computed using the first-differenced series. The smallest value for each model and missing pattern is marked in boldface.
@@
AL (×$\times 10$) 0.56 1.83 3.81 0.46 0.36 0.33 0.40 0.34
(a) Under Missing pattern I
(b) Under Missing pattern II
@@
Next, we turn to the model parameters estimated from the imputed series. Table 2 presents the RMSEs for each model parameter. For DGP 1 (AR) and DGP 2 (ARMA), Kalman smoothing and the scalar filter yield satisfactory estimates among the benchmarks, with TWI offering only limited improvements. For the nonlinear DGP 3 (TAR), the results are drastically different. TWI produces the most accurate parameter estimates and can achieve more than 50% reduction in estimation errors compared to linear interpolation and Kalman smoothing. These estimation results demonstrate TWI’s strength in learning the underlying nonlinear dynamics. For DGP 4 (I(1)), TWI either outperforms or stays on par with the best benchmark, with Kalman filter being a strong initialization.
Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
Missing pattern I
AR $\phi$ 0.06 0.06 0.05 0.03 0.03 0.04 0.03 0.04
ARMA $\phi_{1}$ 0.19 0.47 0.12 0.08 0.17 0.16 0.11 0.11
$\phi_{2}$ 0.46 0.95 0.18 0.10 0.33 0.26 0.16 0.15
TAR $\phi_{1}$ 1.02 1.67 0.82 0.63 0.71 0.58 0.53 0.44
$\phi_{2}$ 0.08 0.07 0.20 0.04 0.05 0.03 0.03 0.02
$\tau$ 0.01 0.75 0.35 0.20 0.05 0.05 0.03 0.03
I(1) $\phi$ 0.96 0.47 0.45 0.68 0.48 0.32 0.30 0.24
NLVAR $\phi_{11}$ 0.19 0.27 0.15 0.03 0.06 0.03 0.05 0.03
$\phi_{12}$ 3.10 3.16 2.87 1.67 1.34 1.02 1.22 0.97
$\phi_{21}$ 0.03 0.07 0.03 0.03 0.03 0.03 0.03 0.03
$\phi_{21}$ 0.18 0.23 0.10 0.04 0.05 0.04 0.04 0.04
Missing pattern II
AR $\phi$ 0.05 0.07 0.04 0.04 0.03 0.02 0.03 0.02
ARMA $\phi_{1}$ 0.04 0.08 0.09 0.08 0.05 0.06 0.09 0.09
$\phi_{2}$ 0.16 0.61 0.11 0.10 0.12 0.11 0.11 0.12
TAR $\phi_{1}$ 0.72 2.44 0.38 0.25 0.25 0.20 0.29 0.21
$\phi_{2}$ 0.09 0.12 0.05 0.03 0.06 0.05 0.02 0.01
$\tau$ 0.01 1.72 0.01 0.01 0.01 0.01 0.01 0.01
I(1) $\phi$ 0.08 1.08 0.11 0.70 0.09 0.10 0.10 0.10
NLVAR $\phi_{11}$ 0.19 0.44 0.08 0.02 0.02 0.01 0.02 0.01
$\phi_{12}$ 2.67 3.51 1.68 1.26 0.61 0.46 0.61 0.43
$\phi_{21}$ 0.02 0.09 0.03 0.03 0.06 0.05 0.04 0.03
$\phi_{21}$ 0.15 0.34 0.05 0.04 0.04 0.04 0.04 0.04
5.2 Multivariate time series
In this subsection, we consider multivariate nonlinear time series models.
DGP 6 (Nonlinear VAR; NLVAR).
$x_{t,1}=\phi_{11}x_{t-1,1}+\phi_{12}\mathfrak{s}(3x_{t-1,2})+0.25\epsilon_{t,1}\quad x_{t,2}=\phi_{21}x_{t-1,1}+\phi_{22}x_{t-1,2}+3\epsilon_{t,2}$
DGP 6 is a VAR model with a nonlinear term. In particular, the impact of $x_{t-1,2}$ on $x_{t,1}$ is nonlinear and is modeled by a sigmoid function. As depicted in Figure 3, TWI successfully captures the underlying nonlinear dynamics, whereas other imputation methods fail to preserve meaningful relationships. This is validated by the results shown in Table 1, where TWI achieves the lowest Wasserstein distance. Next, we assess the effectiveness of the imputed series in estimating the model parameters $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})$, assuming the nonlinear functional form is known. The root mean square estimation errors are documented in Table 2. Compared to the benchmarks, TWI provides considerable error reduction in the parameter associated with the nonlinear effect, $\phi_{12}$. It attains more than 50% reduction in RMSE compared to the scalar filter, the best-performing benchmark.
We also consider a challenging multivariate nonlinear model, which generates the so-called compositional time series. The data are generated from the additive logistic (AL) model of Brunsdon and Smith, (1998). Due to space constraints, we present its simulation results and discussion in Section S3.2 of the supplementary material.
(a) $\hat{w}_{t,1}$ on $\hat{w}_{t-1,1}$
(b) $\hat{w}_{t,2}$ on $\hat{w}_{t-1,1}$
(c) $\hat{w}_{t,1}$ on $\hat{w}_{t-1,2}$
(d) $\hat{w}_{t,2}$ on $\hat{w}_{t-1,2}$